Need some help with absolute value equations. Thanks
[x+2]=[4-x]

Answer by stanbon(48546)

Need some help with absolute value equations. Thanks
[x+2]=[4-x]
---------------
It's a bit of a mess with a simple solution.
A) x+2 = |4-x| or B) x+2 = -|4-x|
------------------------------------
Working with A you get:
4-x = x+2 or 4-x = -(x+2)
2x = 2 or 4-x = -x - 2
x = 1 or contradiction
=============================
Working with B you get:
-(4-x) = x+2 or -(4-x) = -(x+2)
-4+x = x+2 or -4+x = -x-2
-4 = 2 or 2x = 2
contradiction or x =1
===============================
Conclusion: x = 1 is the only solution.
===============================
Note: I graphed it and that answer is true.
===============================================
Cheers,
Stan H.